Find the value of JG. 4х – 10 | 2х + 12 ЈG = 11 ЈG = 18 JG = 22 JG = 34


1
Expert's answer
2021-02-03T04:51:33-0500

in a triangle JGW, WJ = 4x-10, GW = 2x+12, and JG has to find where, all three sides of the triangle are equal.


thus, JG = GW = WJ


now, GW = WJ


2x+12 = 4x-10

2x = 22

x = 11


thus, GW = 2x+12 = (2*11)+12

GW = 34

we know that, JG =GW


therefore, JG = 34